# Differentiable and accelerated spherical harmonic and Wigner transforms

### v1

@@ 1 Introduction
Research efforts in many fields of science and engineering, both in industrial and academic settings, are increasingly considering the analysis of data that live on spherical manifolds, or variants thereof. The diversity in applications is remarkable, ranging from quantum chemistry Ritchie and Kemp (1999); Choi et al. (1999), molecular modelling Boomsma and Frellsen (2017); Kondor et al. (2018), biomedical imaging Tuch (2004); Daducci et al. (2011); McEwen and Price (2019); Goodwin-Allcock et al. (2022), geophysical inverse problems Audet (2011); Simons et al. (2011, 2011); Marignier et al. (2022), general relativity and gravitational waves Thorne (1980); Beyer et al. (2014); Boyle (2016), to the wider cosmos Wallis et al. (2022); Price et al. (2021); Loureiro et al. (2022); Atkins et al. (2023) and beyond. Reflecting the growth of scientific interest, consortiums within these fields have grown in both size and scope, with international missions including, e.g., ESA’s Planck Planck Collaboration I (2020) and Euclid Pires et al. (2020) satellite missions, the Dark Energy Survey (DES; Jeffrey et al., 2021), the Rubin Observatory Legacy Survey of Space and Time (LSST; Ivezić et al., 2019) and soon the Laser Interferometer Space Antenna (LISA; Amaro-Seoane et al., 2017). As such, a substantial global community of researchers have an interest in the development and use of foundational technology for the analysis and modelling of spherical data.
@@
For many of these applications a harmonic analysis of data can be insightful, in some cases critical. Many physical models are best described in harmonic space (often due to the independent physical evolution of different harmonic modes) necessitating spherical harmonic transforms, for example in cosmology (Dodelson, 2003) and seismology (Dahlen and Tromp, 1998). Furthermore, many analyses either rely on spherical harmonic transforms for a direct analysis or as an integral component of other analysis techniques, for example wavelets on the sphere (e.g. McEwen et al., 2007, 2018, 2015) or geometric machine learning on the sphere (e.g. Cohen et al., 2018; Kondor et al., 2018; Esteves et al., 2018; Cobb et al., 2021; McEwen et al., 2022).
@@
Widespread uses of the harmonic analysis of spherical data, combined with increasingly large data volumes, heightens the need for scalable and efficient algorithms, with readily accessible and easy to use software, to compute generalised fast Fourier transforms on both the sphere ${\mathbb{S}^{2}}$ and rotation group ${\text{SO}(3)}$. Modern hardware accelerators, such as GPUs and TPUs, provide a great opportunity to leverage high throughput computing to address the increasing computational challenge. With the growth of differentiable programming opening up many new types of analysis, many applications also require spherical transforms that are differentiable. Machine learning models on the sphere require differentiable transforms so that the models may be trained by gradient-based optimisation algorithms. Emerging physics-enhanced machine learning approaches Karniadakis et al. (2021) also require differentiable physical models, which in many cases themselves require differentiable spherical transforms, for hybrid data-driven and model-based approaches (e.g. Liaudat et al., 2021; Mars et al., 2023; McEwen et al., 2023; Campagne et al., 2023; Piras and Mancini, 2023). While a number of software codes are available to compute spherical harmonic transforms, these do not typically run on hardware accelerators (e.g. GPUs) or support differentiable transforms.
@@
(a) McEwen & Wiaux (MW) McEwen and Wiaux (2011)
(b) Gauss-Legendre (GL) McEwen and Wiaux (2011); Schaeffer (2013)
(c) Driscoll & Healy (DH) Driscoll and Healy (1994)
(d) HEALPix Górski et al. (2005)
Fig. 1: Visualisation of sample positions on the sphere for various sampling schemes. The Driscoll & Healy (DH; Driscoll and Healy, 1994), McEwen & Wiaux (MW; McEwen and Wiaux, 2011), and Gauss-Legendre (GL; McEwen and Wiaux, 2011; Schaeffer, 2013) samplings all admit sampling theorems with exact spherical harmonic transforms, although HEALPix does not. Both MW and GL sampling require $\sim 2L^{2}$ samples (although the MW sampling is slightly more efficient), whereas DH Driscoll and Healy (1994) sampling requires $\sim 4L^{2}$, and HEALPix considers $\sim kL^{2}$ samples, where $k$ typically ranges from $3$ to $12$. The HEALPix scheme has the practical advantage of equal-area pixels and is widely adopted in cosmology. Modern computer vision, machine learning and geophysical applications typically adopt equiangular sampling and thus map most efficiently and straightforwardly onto the MW sampling.
@@ 2.2 Spherical harmonic transforms for spin functions
Spin spherical harmonics may be defined through actions of spin raising operator $\eth$ and lowering operator $\bar{\eth}$ on the scalar spherical harmonics, which are given by Newman and Penrose (1966); Goldberg et al. (1967); Kamionkowski et al. (1997)
@@ 2.3 Wigner transforms for functions on the rotation group
In many settings we encounter functions defined over the space of three-dimensional rotations. The Wigner $D$-functions $D^{\ell}_{mn}:\text{SO}(3)\rightarrow\mathbb{C}$ form an irreducible unitary representation of $\text{SO}(3)$, the group of three-dimensional rotations, for natural $\ell\in\mathbb{N}$ and integers $m,n\in\mathbb{Z}$ such that $|m|,|n|\leq\ell$ Varshalovich et al. (1989). The Wigner $D$-functions are related to the spin spherical harmonics by
@@
where $(\alpha,\beta,\gamma)$ are the Euler angles parameterising $\text{SO}(3)$, with $\alpha\in[0,2\pi)$, $\beta\in[0,\pi]$ and $\gamma\in[0,2\pi)$. We adopt the $zyz$ Euler convention corresponding to the rotation of a physical body in a fixed co-ordinate system about the $z$, $y$ and $z$ axes by $\gamma$, $\beta$ and $\alpha$ respectively. The Wigner $D$-functions may be decomposed in terms of the real Wigner $d$-functions by Varshalovich et al. (1989)
@@
Due to the relationship between the Wigner $D$-functions and the spin spherical harmonics (Eq. 9), Wigner transforms can be expressed as a series of spin spherical harmonic transforms and Fourier transforms McEwen et al. (2015). For completeness, the forward Wigner transform may be expressed as
@@
Note that Eq. 18 corresponds to an inverse spin spherical harmonic transform (of scaled harmonic coefficients) with spin $-n$. These expressions may be leveraged to compute Wigner transforms simply through spin spherical harmonic transforms and standard 1D Fourier transforms McEwen et al. (2015).
@@ 2.4 Sampling schemes versus sampling theorems on the sphere and rotation group
Often we consider the space of bandlimited signals $\mathcal{B}_{L}(\Omega)$, where $\Omega$ represents either the sphere or rotation group, i.e. $\Omega=\{\mathbb{S}^{2},\text{SO}(3)\}$. Bandlimited signals are those whose harmonic coefficients are zero for degrees $\ell$ greater than or equal to the bandlimit $L$. A bandlimited signal on the sphere has $\hat{f}_{\ell m}=0$ for all $\ell\geq L$, while a bandlimited signal on the rotation group has $\hat{f}^{\ell}_{mn}=0$ for all $\ell\geq L$. Consequently, harmonic coefficients for bandlimited signals on the sphere and rotation group live, respectively, in $\mathbb{C}^{L^{2}}$ and $\mathbb{C}^{(4L^{3}-L)/3}$ McEwen et al. (2015); Khalid et al. (2016). For bandlimited signals, summations over degree $\ell$ (cf. Eq. 4 and Eq. 13) may be truncated to $L-1$. In many settings azimuthal bandlimits also arise, such that harmonic coefficients are zero for $m\geq M$ and/or $n\geq N$ (summations over $m$ and $n$ may than also be truncated). Real-world signals may be approximated accurately by bandlimited signals for a suitable bandlimit.
For the spatial signal representation, we distinguish between sampling schemes and theorems. For sampling schemes, the sphere or rotation group may be sampled in any arbitrary manner and harmonic coefficients computed by approximate quadrature. This provides a great deal of flexibility in how sample positions are set. However, this flexibility comes at a cost: for an arbitrary sampling, harmonic transforms can only be computed approximately and in some cases can be quite inaccurate. Furthermore, for an arbitrary sampling the computation of the spherical harmonic transform scales as $\mathcal{O}(L^{4})$. By adopting an isolatitude sampling, where $\varphi$ samples are collected in isolatitudinal rings for each colatitude $\vartheta$, one may apply a separation of variables to reduce computational complexity to $\mathcal{O}(L^{3})$ (e.g. Górski et al., 2005; McEwen and Wiaux, 2011; McEwen et al., 2015, 2013). For this reason isolatitudinal sampling is typically adopted; hence, we focus on isolatitudinal sampling in the remainder of this article. A sampling theorem, in contrast, provides a way to capture all of the information content of a bandlimited signal in a finite set of sampled values. Since all information content of the underlying continuous signal is captured, a sampling theorem implicitly provides a quadrature rule to compute harmonic coefficients exactly. Sampling theorems are thus tied to particular structured samplings of the sphere.
@@ 2.5 Specific isolatitudinal sampling schemes and theorems on the sphere and rotation group
To complete this section, let us briefly discuss the sampling schemes and theorems considered throughout this article and in the associated software implementation. One of the most common spherical sampling schemes is the Hierarchical Equal Area isoLatitude Pixelisation of the sphere, abbreviated HEALPix Górski et al. (2005). [^1] https://healpix.sourceforge.io/ HEALPix has numerous practical advantages, such as a hierarchical representation and equal area pixels. Consequently, it is widely used in many application domains (e.g. Planck Collaboration I, 2020). However, since HEALPix does not exhibit a sampling theorem, spherical harmonic transforms are necessarily approximate. In fact, harmonic transforms can be quite inaccurate with absolute errors of $\sim 10^{-1}$ in some settings Leistedt et al. (2013). For this reason, iterations are usually performed to increase accuracy.
@@
The canonical sampling theorems on the sphere for equiangular samplings are those of Driscoll & Healy (DH; Driscoll and Healy, 1994) and McEwen & Wiaux (MW; McEwen and Wiaux, 2011). Spherical sampling theorems differ significantly to the Euclidean setting. The Nyquist-Shannon sampling theorem for Euclidean spaces results in an identical number of samples in both the spatial and harmonic domains. In contrast, no sampling theorem on the sphere achieves the optimal dimensionality of bandlimited spherical signals in harmonic space, given by $L^{2}$. The MW sampling theorem provides the most efficient equiangular sampling, with $\sim 2L^{2}$ spatial samples, reducing the Nyquist rate on the sphere by a factor of two compared to the DH sampling scheme, which requires $\sim 4L^{2}$ spatial samples. [^2] Note that an equiangular sampling scheme that achieves the optimal number of spatial samples has been developed Khalid et al. (2014). A sampling theorem with exact quadrature is not recovered but high accuracy is nevertheless obtained. Fast algorithms for both the DH and MW sampling theorems have been developed. While a number of algorithms have been introduced for the DH sampling theorem Driscoll and Healy (1994); Healy et al. (2003), some of which have computational complexity of $\mathcal{O}(L^{2}\log L)$, the only variant is that is universally stable is the so-called semi-naive algorithm, with complexity of $\mathcal{O}(L^{3})$. Fast algorithms for the MW sampling theorem also achieve complexity of $\mathcal{O}(L^{3})$ McEwen and Wiaux (2011). While equiangular samplings are the most common structured sampling, for example matching the acquisition format of 360 panoramic photos and videos, spherical sampling theorems based on Gauss-Legendre (GL) quadrature have also been proposed (e.g. Skukowsky, 1986; McEwen and Wiaux, 2011), with samples defined on isolatitudinal rings with $\theta$ specified by the roots of the Legendre polynomials of order $L$. GL sampling exhibits a similar, although slightly greater, number of samples as the MW sampling theorem.
@@
Both the DH and MW sampling theorems have been extended to the rotation group $\text{SO}(3)$ Kostelec and Rockmore (2008); McEwen et al. (2015). Both exhibit fast algorithms with complexity $\mathcal{O}(L^{4})$, while the MW sampling theorem again achieves a reduction in the Nyquist rate on $\text{SO}(3)$ by a factor of two, requiring $\sim 4L^{3}$ spatial samples McEwen et al. (2015) rather than $\sim 8L^{3}$ samples Kostelec and Rockmore (2008). Note that in many practical settings low azimuthal bandlimits arise, reducing a factor of $L$ in both number of samples and complexity to $N\ll L$. The Gauss-Legendre sampling theorem has also been extended to $\text{SO}(3)$ Khalid et al. (2016), requiring a similar (although slightly larger) number of samples as the MW sampling theorem on $\text{SO}(3)$.
@@ 3 Parallelised and differentiable Wigner computation
To compute spin spherical harmonic and Wigner transforms it is necessary to compute the Wigner $D$-functions, or equivalently the real Wigner $d$-functions (Eq. 10). Since we target accelerated computation on modern hardware accelerators, such as GPUs and TPUs that can execute a very large number of threads in parallel, it is desirable to compute Wigner functions in a highly parallelised manner. Furthermore, since we also target differentiable transforms, we must compute Wigner functions in a manner that lends itself to efficient automatic differentiation. Finally, we are interested in use cases that reach very high harmonic degrees $\ell$, hence careful attention must be given to ensure stable computation up to very high $\ell$. In this section we first define explicit expressions for the Wigner functions, before discussing in greater detail the specific requirements that must be met to achieve a high degree of parallel computation and efficient automatic differentiation. We then present a recursive algorithm to compute Wigner functions that meets these requirements, discussing specific considerations to ensure our approach is suitable for differentiable programming frameworks such as JAX Bradbury et al. (2018).

### v2

@@ 1 Introduction
Research efforts in many fields of science and engineering, both in industrial and academic settings, are increasingly considering the analysis of data that live on spherical manifolds, or variants thereof. The diversity in applications is remarkable, ranging from quantum chemistry [1, 2], molecular modelling [3, 4], biomedical imaging [5, 6, 7, 8], geophysical inverse problems [9, 10, 11, 12], general relativity and gravitational waves [13, 14, 15], to the wider cosmos [16, 17, 18, 19] and beyond. Reflecting the growth of scientific interest, consortiums within these fields have grown in both size and scope, with international missions including, e.g., ESA’s Planck [20] and Euclid [21] satellite missions, the Dark Energy Survey [DES; 22], the Rubin Observatory Legacy Survey of Space and Time [LSST; 23] and soon the Laser Interferometer Space Antenna [LISA; 24]. As such, a substantial global community of researchers have an interest in the development and use of foundational technology for the analysis and modelling of spherical data.
@@
For many of these applications a harmonic analysis of data can be insightful, in some cases critical. Many physical models are best described in harmonic space (often due to the independent physical evolution of different harmonic modes) necessitating spherical harmonic transforms, for example in cosmology [25] and seismology [26]. Furthermore, many analyses either rely on spherical harmonic transforms for a direct analysis or as an integral component of other analysis techniques, for example wavelets on the sphere [e.g. 27, 28, 29] or geometric machine learning on the sphere [e.g. 30, 4, 31, 32, 33].
@@
Widespread uses of the harmonic analysis of spherical data, combined with increasingly large data volumes, heightens the need for scalable and efficient algorithms, with readily accessible and easy to use software, to compute generalised fast Fourier transforms on both the sphere ${\mathbb{S}^{2}}$ and rotation group ${\text{SO}(3)}$. Modern hardware accelerators, such as GPUs and TPUs, provide a great opportunity to leverage high throughput computing to address the increasing computational challenge. With the growth of differentiable programming opening up many new types of analysis, many applications also require spherical transforms that are differentiable. Machine learning models on the sphere require differentiable transforms so that the models may be trained by gradient-based optimisation algorithms. Emerging physics-enhanced machine learning approaches [34] also require differentiable physical models, which in many cases themselves require differentiable spherical transforms, for hybrid data-driven and model-based approaches [e.g. 35, 36, 37, 38, 39]. While a number of software codes are available to compute spherical harmonic transforms, these do not typically run on hardware accelerators (e.g. GPUs) or support differentiable transforms.
@@
(a) McEwen & Wiaux (MW) [41]
(b) Gauss-Legendre (GL) [41, 44]
(c) Driscoll & Healy (DH) [40]
(d) HEALPix [43]
Fig. 1: Visualisation of sample positions on the sphere for various sampling schemes. The Driscoll & Healy [DH; 40], McEwen & Wiaux [MW; 41], and Gauss-Legendre [GL; 41, 44] samplings all admit sampling theorems with exact spherical harmonic transforms, although HEALPix does not. Both MW and GL sampling require $\sim 2L^{2}$ samples (although the MW sampling is slightly more efficient), whereas DH [40] sampling requires $\sim 4L^{2}$, and HEALPix considers $\sim kL^{2}$ samples, where $k$ typically ranges from $3$ to $12$. The HEALPix scheme has the practical advantage of equal-area pixels and is widely adopted in cosmology. Modern computer vision, machine learning and geophysical applications typically adopt equiangular sampling and thus map most efficiently and straightforwardly onto the MW sampling.
@@ 2.2 Spherical harmonic transforms for spin functions
Spin spherical harmonics may be defined through actions of spin raising operator $\eth$ and lowering operator $\bar{\eth}$ on the scalar spherical harmonics, which are given by [45, 46, 47]
@@ 2.3 Wigner transforms for functions on the rotation group
In many settings we encounter functions defined over the space of three-dimensional rotations. The Wigner $D$-functions $D^{\ell}_{mn}:\text{SO}(3)\rightarrow\mathbb{C}$ form an irreducible unitary representation of $\text{SO}(3)$, the group of three-dimensional rotations, for natural $\ell\in\mathbb{N}$ and integers $m,n\in\mathbb{Z}$ such that $|m|,|n|\leq\ell$ [48]. The Wigner $D$-functions are related to the spin spherical harmonics by
@@
where $(\alpha,\beta,\gamma)$ are the Euler angles parameterising $\text{SO}(3)$, with $\alpha\in[0,2\pi)$, $\beta\in[0,\pi]$ and $\gamma\in[0,2\pi)$. We adopt the $zyz$ Euler convention corresponding to the rotation of a physical body in a fixed co-ordinate system about the $z$, $y$ and $z$ axes by $\gamma$, $\beta$ and $\alpha$ respectively. The Wigner $D$-functions may be decomposed in terms of the real Wigner $d$-functions by [48]
@@
Due to the relationship between the Wigner $D$-functions and the spin spherical harmonics (Eq. 9), Wigner transforms can be expressed as a series of spin spherical harmonic transforms and Fourier transforms [42]. For completeness, the forward Wigner transform may be expressed as
@@
Note that Eq. 18 corresponds to an inverse spin spherical harmonic transform (of scaled harmonic coefficients) with spin $-n$. These expressions may be leveraged to compute Wigner transforms simply through spin spherical harmonic transforms and standard 1D Fourier transforms [42].
@@ 2.4 Sampling schemes versus sampling theorems on the sphere and rotation group
Often we consider the space of bandlimited signals $\mathcal{B}_{L}(\Omega)$, where $\Omega$ represents either the sphere or rotation group, i.e. $\Omega=\{\mathbb{S}^{2},\text{SO}(3)\}$. Bandlimited signals are those whose harmonic coefficients are zero for degrees $\ell$ greater than or equal to the bandlimit $L$. A bandlimited signal on the sphere has $\hat{f}_{\ell m}=0$ for all $\ell\geq L$, while a bandlimited signal on the rotation group has $\hat{f}^{\ell}_{mn}=0$ for all $\ell\geq L$. Consequently, harmonic coefficients for bandlimited signals on the sphere and rotation group live, respectively, in $\mathbb{C}^{L^{2}}$ and $\mathbb{C}^{(4L^{3}-L)/3}$ [42, 49]. For bandlimited signals, summations over degree $\ell$ (cf. Eq. 4 and Eq. 13) may be truncated to $L-1$. In many settings azimuthal bandlimits also arise, such that harmonic coefficients are zero for $m\geq M$ and/or $n\geq N$ (summations over $m$ and $n$ may than also be truncated). Real-world signals may be approximated accurately by bandlimited signals for a suitable bandlimit.
For the spatial signal representation, we distinguish between sampling schemes and theorems. For sampling schemes, the sphere or rotation group may be sampled in any arbitrary manner and harmonic coefficients computed by approximate quadrature. This provides a great deal of flexibility in how sample positions are set. However, this flexibility comes at a cost: for an arbitrary sampling, harmonic transforms can only be computed approximately and in some cases can be quite inaccurate. Furthermore, for an arbitrary sampling the computation of the spherical harmonic transform scales as $\mathcal{O}(L^{4})$. By adopting an isolatitude sampling, where $\varphi$ samples are collected in isolatitudinal rings for each colatitude $\vartheta$, one may apply a separation of variables to reduce computational complexity to $\mathcal{O}(L^{3})$ [e.g. 43, 41, 42, 50]. For this reason isolatitudinal sampling is typically adopted; hence, we focus on isolatitudinal sampling in the remainder of this article. A sampling theorem, in contrast, provides a way to capture all of the information content of a bandlimited signal in a finite set of sampled values. Since all information content of the underlying continuous signal is captured, a sampling theorem implicitly provides a quadrature rule to compute harmonic coefficients exactly. Sampling theorems are thus tied to particular structured samplings of the sphere.
@@ 2.5 Specific isolatitudinal sampling schemes and theorems on the sphere and rotation group
To complete this section, let us briefly discuss the sampling schemes and theorems considered throughout this article and in the associated software implementation. One of the most common spherical sampling schemes is the Hierarchical Equal Area isoLatitude Pixelisation of the sphere, abbreviated HEALPix [43]. [^1] https://healpix.sourceforge.io/ HEALPix has numerous practical advantages, such as a hierarchical representation and equal area pixels. Consequently, it is widely used in many application domains [e.g. 20]. However, since HEALPix does not exhibit a sampling theorem, spherical harmonic transforms are necessarily approximate. In fact, harmonic transforms can be quite inaccurate with absolute errors of $\sim 10^{-1}$ in some settings [51]. For this reason, iterations are usually performed to increase accuracy.
@@
The canonical sampling theorems on the sphere for equiangular samplings are those of Driscoll & Healy [DH; 40] and McEwen & Wiaux [MW; 41]. Spherical sampling theorems differ significantly to the Euclidean setting. The Nyquist-Shannon sampling theorem for Euclidean spaces results in an identical number of samples in both the spatial and harmonic domains. In contrast, no sampling theorem on the sphere achieves the optimal dimensionality of bandlimited spherical signals in harmonic space, given by $L^{2}$. The MW sampling theorem provides the most efficient equiangular sampling, with $\sim 2L^{2}$ spatial samples, reducing the Nyquist rate on the sphere by a factor of two compared to the DH sampling scheme, which requires $\sim 4L^{2}$ spatial samples. [^2] Note that an equiangular sampling scheme that achieves the optimal number of spatial samples has been developed [52]. A sampling theorem with exact quadrature is not recovered but high accuracy is nevertheless obtained. Fast algorithms for both the DH and MW sampling theorems have been developed. While a number of algorithms have been introduced for the DH sampling theorem [40, 53], some of which have computational complexity of $\mathcal{O}(L^{2}\log L)$, the only variant is that is universally stable is the so-called semi-naive algorithm, with complexity of $\mathcal{O}(L^{3})$. Fast algorithms for the MW sampling theorem also achieve complexity of $\mathcal{O}(L^{3})$ [41]. While equiangular samplings are the most common structured sampling, for example matching the acquisition format of 360 panoramic photos and videos, spherical sampling theorems based on Gauss-Legendre (GL) quadrature have also been proposed [e.g. 54, 41], with samples defined on isolatitudinal rings with $\theta$ specified by the roots of the Legendre polynomials of order $L$. GL sampling exhibits a similar, although slightly greater, number of samples as the MW sampling theorem.
@@
Both the DH and MW sampling theorems have been extended to the rotation group $\text{SO}(3)$ [55, 42]. Both exhibit fast algorithms with complexity $\mathcal{O}(L^{4})$, while the MW sampling theorem again achieves a reduction in the Nyquist rate on $\text{SO}(3)$ by a factor of two, requiring $\sim 4L^{3}$ spatial samples [42] rather than $\sim 8L^{3}$ samples [55]. Note that in many practical settings low azimuthal bandlimits arise, reducing a factor of $L$ in both number of samples and complexity to $N\ll L$. The Gauss-Legendre sampling theorem has also been extended to $\text{SO}(3)$ [49], requiring a similar (although slightly larger) number of samples as the MW sampling theorem on $\text{SO}(3)$.
@@ 3 Parallelised and differentiable Wigner computation
To compute spin spherical harmonic and Wigner transforms it is necessary to compute the Wigner $D$-functions, or equivalently the real Wigner $d$-functions (Eq. 10). Since we target accelerated computation on modern hardware accelerators, such as GPUs and TPUs that can execute a very large number of threads in parallel, it is desirable to compute Wigner functions in a highly parallelised manner. Furthermore, since we also target differentiable transforms, we must compute Wigner functions in a manner that lends itself to efficient automatic differentiation. Finally, we are interested in use cases that reach very high harmonic degrees $\ell$, hence careful attention must be given to ensure stable computation up to very high $\ell$. In this section we first define explicit expressions for the Wigner functions, before discussing in greater detail the specific requirements that must be met to achieve a high degree of parallel computation and efficient automatic differentiation. We then present a recursive algorithm to compute Wigner functions that meets these requirements, discussing specific considerations to ensure our approach is suitable for differentiable programming frameworks such as JAX [56].